# The Power of Training: How Different Neural Network Setups Influence the Energy Demand

### 03 Jan 2024 v1

@@ Abstract
This work examines the effects of variations in machine learning training regimes and learning paradigms on the corresponding energy consumption. While increasing data availability and innovation in high-performance hardware fuels the training of sophisticated models, it also supports the fading perception of energy consumption and carbon emission. Therefore, the goal of this work is to create awareness about the energy impact of general training parameters and processes, from learning rate over batch size to knowledge transfer. Multiple setups with different hyperparameter initializations are evaluated on two different hardware configurations to obtain meaningful results. Experiments on pretraining and multitask training are conducted on top of the baseline results to determine their potential towards sustainable machine learning.
Introduction
The connection between the complexity of a task being solved by a deep neural network and the accompanying energy requirement from powerful hardware to train and deploy such models is a well-known issue of machine learning algorithms. Improving availability and access to more data, which forms the basis for training complex models, supports the trend towards enhanced hardware power and energy consumption further. Next to the increase in computational demands, there is a growing awareness of a less visible but increasingly significant aspect: the carbon footprint for training machine learning models. (Henderson et al. 2020; Patterson et al. 2022)
Multiple projects like carbontracker (Anthony, Kanding, and Selvan 2020) or eco2AI (Budennyy et al. 2022) apply to this problem statement, aiming to track the energy consumption and the related carbon footprint for training a model by linking into the hardware power utilization logging during the model training. Although these tools can monitor energy consumption, they are unable to compare the efficiency of the training process as well as suggesting recommendations for improvements.
@@ Introduction
Even though the innovation in terms of efficiency and sustainability of machine learning algorithms lags behind quantitative growth of hardware resources, small improvements in the low percentage range already help to reduce the carbon footprint if such methods are scaled properly within the landscape of machine learning applications. (Patterson et al. 2022)
Related Works
@@
Carbon tracking software tools are usually designed to capture as much power information about the system as possible by building an additional layer between the system’s hardware configuration and the user’s model training process. Generally speaking, the power consumption of the GPU is the largest part of the training process as it performs the core work with the parallel processing of mathematical tasks. Following that, usually, the CPU is the second largest consumer of power.
@@
There is a growing list of software available with most of them utilizing the same approach to gather the power consumption data from the hardware manufacturers’ utility logging. The energy consumption is then calculated from the current power consumption and the polling time interval set in the software. Works like Carbontracker (Anthony, Kanding, and Selvan 2020), eco2Ai (Budennyy et al. 2022) and Green Algorithms (Lannelongue, Grealey, and Inouye 2021) utilize this approach to gather data from CPU, GPU, and even the RAM if supported. Nevertheless, there is a lack of tracking the full systems’ energy consumption. High-performance setups may consume more power than present tools can track, for instance, if the cooling, which can account for a non-negligible proportion, is not taken into account. Therefore, software like Carbontracker (Anthony, Kanding, and Selvan 2020) multiplies its results with an efficiency constant of 1.55 to incorporate untracked secondary power needs and efficiency losses.
@@
With an extended focus on user experience, projects like Cloud Carbon (cloud-carbon footprint 2023) or CodeCarbon (Code-Carbon 2023) extend the gathered knowledge and present it in analytic-based dashboards. Based on the calculated energy consumption and the user’s location, the average local energy mix from fossil and renewable energy sources is utilized to estimate the carbon emissions in kilograms or even tons. (Lacoste et al. 2019) On top of that, since the carbon emissions are difficult to visualize or imagine, the conversion into kilometers driven by car, flights with a plane, or the number of phones charged is a standard practice to make the user aware of the generated carbon emissions amount.
Methodology
Experiment Setup
@@
In order to provide a meaningful correlation between the training setup of a model and its power consumption, we chose two application scenarios that determine the structure of this work: computer vision and sensor-based activity recognition. We used two commonly used benchmark datasets from two disciplines: for a computer vision task we used the CelebA dataset (Liu et al. 2015) and for the sensor-based activity recognition, we utilized the PAMAP2 dataset (Reiss and Stricker 2012) as a basis.
We run each experiment two times on two different hardware configurations to reduce hardware-specific implications from our results. The already introduced Carbontracker (Anthony, Kanding, and Selvan 2020) is used to track the GPU power consumption on a workstation powered by an Nvidia RTX 6000 Ada with 48GB VRAM and AMD Ryzen 9 7950X CPU. A 16-inch Apple Macbook Pro with M1 Max chip with 32-core GPU and 64GB unified memory was used as a second device. The Nvidia GPU workstation runs on CUDA 12.3 and the Apple device runs on metal performance shader (MPS) for neural network training acceleration with pytorch. We implemented a custom tracker based on the powermetrics tool of Mac OS to generate the same data structure as the Carbontracker. Due to Carbontracker’s limitations on tracking CPU power consumption for Intel CPUs only, we distinguished the gathering of GPU and CPU power consumption on the Apple setup as well. As a result, both systems were set up to gather the power consumption of the utilized GPU during the training process in milliwatts precision with a sampling rate of one second. Additionally, the trackers stored the corresponding timestamp and the current epoch to properly evaluate the data. The efficiency constant multiplication was disabled for out testings to achieve comparable measurements.
@@
Throughout the whole experiment and different hardware setups, we kept the same model and training script versions. Moreover, the changed parameters across training runs were induced as external arguments. For training, we used early stopping techniques based on validation loss for the encoder and the validation accuracy for the classifier as a standard metric.
Training regime
Our test procedure can be separated into two tasks which we applied to the two application scenarios, the training regime and the learning paradigm.
A common problem when training a model is the proper initialization of hyperparameters that fit the task or problem. The two most prominent ones are the batch size and the learning rate (He, Liu, and Tao 2019). To determine the influence of hyperparameters, we created an encoder plus classifier architecture for the CelebA to classify the attributes and identities. The encoder is ResNet18 from the TIMM python package with pre-trained ImageNet weights. For the PAMAP2, we build a similar convolutional classifier architecture inspired by (Suh, Rey, and Lukowicz 2023) to classify the activities. For the training regime, we trained the models with variations in batch size and learning rate, resulting in a total of 16 different setups as shown in table 1.
@@
Table 1: Training Regime Hyperparameter Setup List
Learning Paradigm
@@
With the PAMAP2 dataset, we evaluate the pretraining learning paradigm, shown in figure 1. We train an autoencoder and use its encoder as a pre-trained part to train a classifier on its latent features and compare it to the training of the encoder and classifier architecture sequentially. This practice is typically used to first force the feature encoder to learn information that can be used to reconstruct the input, thus the features are unique to different characteristics of the input data. This could reduce the likelihood of over-fitting on the ground truth labels of the training data and thus provide a better generalized model. We also inspect the differences between freezing and unfreezing the pre-trained encoder weights during the usage for generating the latent features.
@@
With the CelebA dataset, we can evaluate multi-task learning (Yin and Liu 2017), since the datasets have different aspects of ground truth for the same input (identity and facial attributes). Opposed to training a complete black-box-like neural network for each task, multi-task learning shares the feature encoder with multiple downstream tasks (e.g. different classifiers). The procedure is visualized in figure 2 a and b for the single training setup and the multitask setup in c by connecting the two classifiers via the weight coefficient $\alpha$ within the loss function.
@@
Training Regime Results
Learning Rate
Batch Size 0.1 0.01 0.001 0.0001
64 109.63 110.42 107.50 108.37
256 42.34 40.20 40.19 38.41
1024 30.84 30.97 30.87 30.81
4096 38.57 37.13 37.44 38.22
(a) PAMAP2 on CUDA (mWh)
Learning Rate
Batch Size 0.1 0.01 0.001 0.0001
64 38.40 39.25 38.73 38.10
256 39.03 40.19 39.91 39.34
1024 43.52 43.98 44.26 42.98
4096 43.00 43.97 44.01 43.09
(b) CelebA on CUDA (mWh)
Learning Rate
Batch Size 0.1 0.01 0.001 0.0001
64 15.40 15.63 15.96 15.66
256 16.83 16.79 16.76 25.35
1024 25.31 25.35 25.45 25.57
4096 27.33 27.17 27.11 27.24
(c) PAMAP2 on MPS (mWh)
Learning Rate
Batch Size 0.1 0.01 0.001 0.0001
64 22.81 22.78 22.75 22.79
256 22.29 22.28 22.31 22.28
1024 22.37 22.34 22.25 22.39
4096 22.39 22.29 22.28 22.37
(d) CelebA on MPS (mWh)
Table 2: Average Energy per Epoch (mWh) across all setup combinations.
(a) Sensor-based activity recognition (PAMAP2) on CUDA
(b) Sensor-based activity recognition (PAMAP2) on MPS
(c) Identification from images (CelebA) on CUDA
(d) Identification from images (CelebA) on CUDA
@@
Figure 3: Average consumption of each setup across all epochs by using the worst performing setup as baseline.
@@
For testing the training regime on the two application scenarios, the standard versions of their tasks are utilized by excluding the advanced setups so far. For PAMAP2, this is the training of an encoder and a decoder in one run without pretraining, whereas for the CelebA scenario, a single classifier is trained to predict the identities only. Each scenario is trained with 16 different hyperparameter combinations and run two times on the two different hardware setups.
@@
The average energy that was demanded during the training is shown in table 2, calculated through the introduced equation 1. As a general rule, the level of energy used can be linked to the efficiency and utilization of the GPU. Lower values mean less energy consumed for training one epoch, which results in improved GPU efficiency.
For each matrix shown in table 2, the batch size choice has the most impact on the energy consumption. There needs to be a balance between the dataset size and the GPU capabilities. On CUDA (matrix 1(a) and 1(b)), the PAMAP2 dataset performs best for bigger batch sizes due to its slimmer data structure size whereas the bigger image dataset CelebA performs best for smaller batches. This is mainly ruled by the GPU utilization, but influenced by the memory bandwidth limitations as well. Similar results can be found for the same runs on MPS (matrix 1(c) and 1(d)), keeping in mind the less powerful system. Therefore, the best batch size choice moves towards the smaller batch sizes for the PAMAP2 scenario and even stays constant for the bigger CelebA dataset due to the system resource limitations.
@@
The choice of learning rate does not have a recognizable influence in this analysis. There is little to no change in the energy per epoch across the four learning rate settings. The reason for that is the independence of the training duration, namely the number of epochs until the model converges and stops the training process. Influences through the learning rate change can therefore be seen in figure 3. Across the 16 setups, grouped into the 4x4 setup alignment through the dashed lines as introduced in table 1, the energy consumption in percentage is visualized, assuming the worst run to be at 100%. We can derive a slight trend of learning rate settings towards 0.01 and 0.001, mainly for the bigger batch sizes. This can be due to the issue of learning rate 0.1 being set too large to converge properly whereas learning rate of 0.0001 increases the number of epochs needed until convergence. Additionally, the batch size seems to have a slight influence on the total energy consumption as well. For instance in figure 2(b) and 2(c), the combination of a small batch size of 64 and the biggest learning rate of 0.1 outperforms the other setups using bigger batch sizes.
Overall, the range between the worst and best-performing setup is wide, with the best model usually consuming only around 20% of the worst model. As a guideline to achieve this much of an improvement, the energy per epoch and the number of epochs necessary need both to be minimized.
Pretraining Learning Paradigm Results

### 15 Mar 2024 v2

@@ Abstract
This work offers a heuristic evaluation of the effects of variations in machine learning training regimes and learning paradigms on the energy consumption of computing, especially HPC hardware with a life-cycle aware perspective. While increasing data availability and innovation in high-performance hardware fuels the training of sophisticated models, it also fosters the fading perception of energy consumption and carbon emission. Therefore, the goal of this work is to raise awareness about the energy impact of general training parameters and processes, from learning rate over batch size to knowledge transfer. Multiple setups with different hyperparameter configurations are evaluated on three different hardware systems. Among many results, we have found out that even with the same model and hardware to reach the same accuracy, improperly set training hyperparameters consume up to 5 times the energy of the optimal setup. We also extensively examined the energy-saving benefits of learning paradigms including recycling knowledge through pretraining and sharing knowledge through multitask training.
Keywords:
Energy Demand Hyperparameter HPC Monitoring
1 Introduction
With the wake of artificial intelligence (AI), the high-performance computing (HPC) demand related to machine learning has been drastically elevated, together with the energy consumption in the sector as a whole. While there have been many works optimizing the energy footprint of machine learning from the software model architecture and hardware accelerator aspects, there has been little effort in examining the differences in training regimes and learning paradigms, especially considering the lifecycle of machine learning models. The connection between the complexity of a task being solved by a deep neural network (DNN) and the accompanying energy requirement from powerful hardware to train and deploy such models is a well-known issue of machine learning algorithms. Improving availability and access to more data, which forms the basis for training complex models, supports the trend towards enhanced hardware power and energy consumption further. Next to the increase in computational demands, there is a growing awareness of a less visible but increasingly significant aspect: the carbon footprint for training machine learning models. [11, 18]
Multiple projects like carbontracker [1] or eco2AI [4] apply to this problem statement, aiming to track the energy consumption and the related carbon footprint for training a model by linking into the hardware power utilization logging during the model training. Although these tools can monitor energy consumption, they are unable to compare the efficiency of the training process as well as suggesting recommendations for improvements.
@@ 1 Introduction
Even though the innovation in terms of efficiency and sustainability of machine learning algorithms lags behind the quantitative growth of hardware resources, small improvements in the low percentage range already help to reduce the carbon footprint if such methods are scaled properly within the landscape of machine learning applications. [18]
In short summary, we make the following contributions:
  1. 1.
    We examine the energy consumption of often overlooked nuance during DNN training, specifically among training setup hyperparameters and learning paradigms with longer lifecycle considerations. A major factor differentiating our method is that we use the prediction accuracy as a performance target, resulting in total energy per development cycle, additionally to the pure energy per epoch metrics like most related works.
  2. 2.
    Our findings show, with the same DNN model, HPC hardware, and accuracy-regulated tasks, improperly configured batch size and learning rate could cause 5 times the energy consumption of the optimal configurations.
  3. 3.
    By re-using part of a DNN (the feature encoder) in different downstream classification tasks, the practice will bring a break-even point compared to training a new model in total for every downstream task. And this break-even point is associated with the training hyperparameter setups in Contribution No. 2.
  4. 4.
    By sharing parts of a DNN during the same training cycle of multiple tasks (multi-task training), the energy consumption per cycle is significantly less than the total of training for different tasks in isolation.
The paper is structured as follows: section 2 reviews the relevant state-of-the-art works to track energy consumption and optimize the hyperparameter space; section 3 introduces the methodology of conducted experiments from classic training regime to advanced learning paradigms followed by the corresponding results for the classic training in section 4, the pretraining scenario in section 5 and multitask scenario in section 6. After additional comparison of the HPC results to high-performance consumer hardware in section 7, we conclude with a discussion on limitations and future work in section 8.
2 Related Works
Due to lacking practicability and the complexity of tasks, precise energy consumption tracking is still not commonly applied in the field of AI and Machine Learning [9]. However, the rising demand for energy must be counteracted to minimize the environmental footprint and economic expenses, for instance when hosting large server and HPC farms, in order to operate sustainably in the future [17]. In the past focus lay on inference optimization for edge devices, microcontrollers, and field-programmable gate arrays (FPGAs), since the energy demands of training machine learning models have received less attention due to smaller implications [6]. Nevertheless, with the rise of Large Language Models (LLM) and other complex architectures, the computational demands for training such models have a significantly larger share of the lifecycle energy requirement [20, 16]. This in turn requires a precise consideration of possible adjustment screws to counteract these trends, whereas this work focuses on the exploration and tuning of hyperparameters [10]. To the best of our knowledge, this work is the first to investigate the selection of popular hyperparameter spaces with a hardware-based focus on quantizing the energy demand instead of purely aiming towards optimal model performance.
2.1 Energy Tracking
@@
Energy tracking software tools are usually designed to capture as much power information about the system as possible by building an additional layer between the system’s hardware configuration and the user’s model training process. Generally speaking, the power consumption of the Graphics Processing Unit (GPU) is the largest part of the training process as it performs the core work with the parallel processing of mathematical tasks. Following that, usually, the Central Processing Unit (CPU) is the second largest consumer of power.
@@
There is a growing list of software available with most of them utilizing the same approach to gather the power consumption data from the hardware manufacturers’ utility logging. The energy consumption is then calculated from the current power consumption and the polling time interval set in the software. Works like Carbontracker [1], eco2Ai [4] and Green Algorithms [13] utilize this approach to gather data from CPU, GPU, and even the memory if supported. Nevertheless, there is still a lack of tracking the full systems’ energy consumption. High-performance setups may consume more power than present tools can track, for instance, if the cooling, which can account for a non-negligible proportion, is not taken into account. Therefore, software like Carbontracker [1] multiplies its results with an efficiency constant of 1.55 to incorporate untracked secondary power needs and efficiency losses.
@@
With an extended focus on user experience, projects like Cloud Carbon [7] or CodeCarbon [5] extend the gathered knowledge and present it in analytic-based dashboards. Based on the calculated energy consumption and the user’s location, the average local energy mix from fossil and renewable energy sources is utilized to estimate the carbon emissions in kilograms or even tons [12]. On top of that, since the carbon emissions are difficult to visualize or imagine, the conversion into kilometers driven by car, flights with a plane, or the number of phones charged is a standard practice to make the user aware of the generated carbon emissions amount.
2.2 Hyperparameter Selection
Selecting the right hyperparameters is a common problem in thoroughly training the machine learning model to properly solve the targeted task. It is still a standard practice to explore the hyperparameter space manually by launching multiple training runs and wasting energy on training insufficient models [22].
Many techniques evolved over time to solve this issue, starting from classical methods like grid search or random search, towards more advanced techniques like evolutionary, bayesian, pruning or reinforcement-learning based optimization strategies [3]. In many cases, there is no clear identification of the best-performing strategy possible because the efficiency of the algorithm depends on the respective machine learning problem to solve and the user’s preferences. Nevertheless, there is a trend of different hyperparameter optimization categories being fused to elevate the performance compared to the traditional algorithms [2].
All strategies together share a downside: They purely focus on maximizing the model performance without incorporating the energy consumption [8]. This oversight leads to an enhanced waste of energy and usually leads to training and discarding models that don’t meet the required performance [25]. Therefore, there is a need to investigate the correlation between hyperparameter selection and energy demand in order to develop strategies that may strive for the best performance while maintaining moderate energy demands.
3 Methodology
3.1 Experiment Setup
@@
In order to provide a meaningful correlation between the training setup of a model and its power consumption, we chose two application scenarios that determine the structure of this work: computer vision and sensor-based activity recognition. We used two commonly used benchmark datasets from two disciplines: for a computer vision task we used the CelebA dataset [14] and for the sensor-based activity recognition, we utilized the PAMAP2 dataset [19] as a basis.
To reduce hardware-specific implications from our results and cross-compare them in a meaningful way, we run each experiment two times on three different hardware configurations on a SLURM-based HPC cluster [24]. Due to the job-based nature of the HPC, the experiments run in clean, isolated Linux environments running on CUDA 12.3. We settled three different Nvidia GPUs, namely the H100-HBM3 with 80GB, the A100-SXM4 with 80GB, and the GeForce RTX 3090.
The already introduced Carbontracker [1] is utilized to track the GPU power and energy consumption of the desired GPUs. Due to Carbontracker’s restriction on tracking CPU power consumption for Intel CPUs only and a lack of an adequate framework to track the power consumption of AMD CPUs, this work solely focuses on GPU power consumption. The tracker is modified to additionally store the corresponding timestamp and the current epoch to properly link the gathered data with the model’s training process. The efficiency constant multiplication was disabled throughout our tests to achieve the raw GPU measurements.
To further classify the HPC results on common prosumer hardware, the same experimental scenarios were carried out again on a workstation equipped with an Nvidia RTX 6000 Ada with 48GB VRAM and a 16-inch Apple Macbook Pro with M1 Max chip (32-core GPU and 64GB unified memory). The findings for those hardware are separately discussed in Section 7.
@@
Throughout all experiments and the varying hardware setups, we kept the same model and training script versions. Moreover, the changed parameters across training runs were induced as external arguments. For training, we used early stopping techniques based on validation loss for the encoder and the validation accuracy for the classifier as a standard metric.
3.2 Training Regime
Our test procedure can be separated into two tasks which we applied to the two application scenarios, the training regime and the learning paradigm. A common problem when training a model is the proper initialization of hyperparameters that fit the task or problem. The two most prominent ones are the batch size and the learning rate [10]. To determine the influence of hyperparameters, we created an encoder plus classifier architecture for the CelebA to classify the attributes and identities. The encoder is ResNet18 from the TIMM python package with pre-trained ImageNet weights and the classifier is based on classic convolution with a size of 20520 parameters. For the PAMAP2, we build a similar convolutional classifier architecture inspired by [21] with a size of 141293 parameters to classify the activities. For the training regime, we trained the models with variations in batch size and learning rate, resulting in a total of 16 different setups as shown in Table 1. We kept the hyperparameters constant for each training run without any further decay or adaption. All experiments were conducted with threshold early stopping as a regularization in order to still maintain proper model performance. We set the threshold to an accuracy of 0.8 from which the regularization was started. It was set based on experience since such classifiers trained on CelebA and PAMAP2 usually range slightly above this performance. This strategy ensures that all setups can exceed the threshold and further eliminates potential outliers due to early termination with poor accuracy.
Table 1: Training Regime Hyperparameter Setup List
@@ 3.2 Training Regime
3.3 Learning Paradigm
@@
With the PAMAP2 dataset, we evaluate the pretraining learning paradigm, shown in Figure 1. We train an autoencoder and use its encoder as a pre-trained part to train a classifier on its latent features and compare it to the training of the encoder and classifier architecture sequentially. This practice is typically used to first force the feature encoder to learn information that can be used to reconstruct the input, thus the features are unique to different characteristics of the input data. This could reduce the likelihood of over-fitting on the ground truth labels of the training data and thus provide a better-generalized model. We also inspect the differences between freezing and unfreezing the pre-trained encoder weights during the usage for generating the latent features.
@@
With the CelebA dataset, we can evaluate multi-task learning [23], since the datasets have different aspects of ground truth for the same input (identity and facial attributes). Opposed to training a complete black-box-like neural network for each task, multi-task learning shares the feature encoder with multiple downstream tasks (e.g. different classifiers). The procedure is visualized in Figure 2 a and b for the single training setup and the multitask setup in c by connecting the two classifiers via the weight coefficient $\alpha$ within the loss function.
@@
4 Training Regime Results
(a) CelebA on H100
(b) CelebA on A100
(c) CelebA on RTX 3090
(d) PAMAP2 on H100
(e) PAMAP2 on A100
(f) PAMAP2 on RTX 3090
Figure 3: Energy per epoch across the different application scenarios with varying Batch Size and Learning Rate.
(a) CelebA on Nvidia H100
(b) PAMAP2 on Nvidia H100
(c) CelebA on Nvidia A100
(d) PAMAP2 on Nvidia A100
(e) CelebA on Nvidia RTX 3090
(f) PAMAP2 on Nvidia RTX 3090
@@ 4 Training Regime Results
Figure 4: Average consumption across all epochs for each application scenario and test setup.
@@
For testing the training regime on the two application scenarios, the standard versions of their tasks are utilized by excluding the advanced setups so far. For PAMAP2, this is the training of an encoder and a classifier in one run without pretraining, whereas, for the CelebA scenario, a single classifier is trained to predict the identities only. Each scenario is trained with 16 different hyperparameter combinations and run two times on the three different HPC hardware setups.
@@
The average energy that was demanded during the training is shown in Figure 3, calculated through the introduced equation Equation 1. As a general rule, the level of energy used can be linked to the efficiency and utilization of the GPU. Within each matrix, lower values mean less energy consumed for training one epoch, which results in improved GPU efficiency.
For each matrix shown in table Figure 3, the batch size choice has the most impact on the energy consumption. Since the batch size determines the level of operations that can be executed in parallel, effects in the energy per epoch mainly originate from the utilization and the efficiency window in which the GPU is operated. There needs to be a balance between the dataset size and the GPU capabilities in order to operate it efficiently. Cross-comparing the matrices across the two application scenarios highlights the importance of proper hyperparameter utilization. Due to the large size of the CelebA image dataset with each data point representing a 78×218 pixel-sized image, the smallest batch size can be considered the best selection to operate the GPU efficiently On the other hand, the tremendously smaller size of the PAMAP2 results in in-efficient utilization of the GPU for small batch sizes. Therefore, larger batch sizes run more efficiently for the PAMAP2 scenario since less energy is wasted for GPU idling.
@@
Compared to the batch size, the choice of learning rate does not have a recognizable influence in this analysis. Apart from small deviations, which may occur due to the software-based measurement procedure, we cannot derive a trend for the energy per epoch across the four learning rate selections. The reason for that is the independence of the training duration, namely the number of epochs until the model converges and stops the training process. As mentioned before, the early stopping was set to start after reaching the threshold accuracy of 0.8 with patience of 40 epochs. Influences through the learning rate change can therefore be seen in figure Figure 4. Across the 16 setups, grouped into the 4x4 setup alignment through the dashed lines as introduced in Table 1, the total energy consumption for the full training is visualized. We can derive a slight trend of learning rate settings towards 0.01 and 0.001, due to the issue of learning rate 0.1 being set too large to converge properly whereas a learning rate of 0.0001 increases the number of epochs needed until convergence. Further, the total consumption is still ruled by the batch size choice, especially in the PAMAP2 scenario with greater deviations in energy per epoch.
Overall, the range between the worst and best-performing setup is wide, with the best case usually consuming less than 20% of the worst-case setup. As a guideline to achieve this much of an improvement, the energy per epoch and necessary number of epochs both need to be minimized. Comparing our two application scenarios, especially when checking the CelebA results in Figure 4, there can not be derived a general rule for thoroughly choosing the best hyperparameters. Considering the extension with less prominent hyperparameters, for instance These findings therefore provide a motivation for further research to explore more dynamic and thorough hyperparameter selection approaches.
5 Pretraining Learning Paradigm Results